Prev Next

SW&IT Meeting Minutes 2020-05-26

Attendees: Carl, David, Igor, Matthieu, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20200518
LR20200518 - Bec, Matthieu 40.0
LR20200518 - Holmberg, Carl 24.0
LR20200518 - Edmeades, David 40.0
LR20200518 - Sola, Igor 40.0
LR20200518 - Hooper, Stephen 80.0
LR20200518 - Zhang, Yang 40.0
LR20200518 - Zhang, Xianyu 0.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

  • telescope / instruments subsystems restart

Team reports

PK

MB

  • Last week
    • OC, Management
    • Technical Requirements for a Remote Observing Facility - draft v0.2
    • catch-up ALTA undo after restart overlooked earlier
    • AO PB/SOUL/ARGOS/LN, draft presentations v0.5
    • screen tests for ROWG in 507/580/584
    • SHARK-VIS updated RFA + input from DE for IT
  • This week
    • DMS
      • websockets, AdSec, etc.
      • LBC stiched png preview
    • cobbler / kickstart, ims f30

DE

  • Luci server spec
  • Cisco training purchase sorted, will start this week
  • Avtech sensors on summit (PG cabinet L4)
  • LeRoy blocked for time Wednesday to work on sensors
  • Re-followup on Crucial SSD; they had not done what they promised before. I am assured that it will happen RSN
  • Notice that Dan was able to register a copy of Inventor, so that's good re: Autodesk licensing
  • Veeam quote requested, waiting on CDW for quote.
  • Josh VPN setup/learning
    • He didn't understand what a VPN does; perhaps we can write a short document with instructions and a basic tutorial on what it is and how it works.
  • Jim MS problems. I think they're solved, but I never get a coherent followup from him, just a cessation of communication.
  • Track Jim printer so I can know when the storm is coming.
  • Modified backup scripts to handle changing DNS environment
  • Possible hiccup re: SMB mounts using fqdn that are set to be depricated. Sometimes this is necessary due to braindead Windows limitation on credentials.

YZ

  • Previous:
    • TMS information review/sketch/meeting.
    • AO-Cravat experiments/meeting organization.
    • TCS/IIF documents rework.
    • Safety training.
  • Next:
    • Some additional AO-Cravat work before the meeting.
    • TMS meeting, revise sketch accordingly.
    • Find some time to continue previous Python projects (MODS scripts and WCS infrastructure).
    • Olga's MODS UI request??
    • Steve Ertel's LBTO github access.
    • Hopefully finish safety training.

SH

  • mountain power up (went pretty well)
  • laptops over to Safford this weekend
  • DNS in Tucson
  • Moved AD to new home
  • LDAP. Google accounts, account cleanup.
  • move at least three developer machines from vm2.tucson
  • ROWG setup in 530 and 538
    • return of DOA ROWG monitor (~20 minutes mainly getting packing material), but just be aware it has happened.
  • wiki.lbto.org page restore

next:

  • obs1 Tucson
  • New vpn.lbto.org
    • setup
    • purchase SSL cert for vpn.mountain.lbto.org
    • have lbto.org point to our authoritative DNS.
    • send out e-mail describing new log on
    • flip DNS
  • RPMs need some work
  • iLocater aggregate port
  • changes to radar image

XZ

  • Last week
    • Finished merging UAO to SOUL
    • Test the merged AO versions on the VM simulator
    • Troubleshooting of SX ASM issue
    • Meetings
  • This week
    • Troubleshooting of SX ASM issue
    • Working on PB action items
    • AdSec restart

CH

Last Week:
  • OA Control: Install DT dev IDE, document existing PMAC code.
  • Safety training
This Week:
  • DMS: commit new sensor handling code to branch.
  • OA Control: document existing PMAC code, review with John Morris, Dan Cox.
  • Argos: get RPM build working.

IS

Last week:
  • LUCI
    • supported Dave and James in the LUCI1 detector testing.
    • Helping Jenny to set up testing of LUCI sym in the obs machines.
    • started looking at #IT7952 an RTD/Aladin GUI issue.
    • started LUCI WebIO device configuration.
  • OT
    • Release last friday
This week:
  • TRAINING
  • LUCI
    • work on #IT7952
    • finish configuration of LUCI WebIO
    • revert to old GEIRS in LUCI2 osa-luci.sim
  • PIT
    • If I find some time, work on the PIT admin tool backend.
  • OT
    • Move QA backend to pit.lbto.org and update release pages.
Next week
  • I will be working from California in June
* Old laptop ready but if it is not inmediately needed I rather have it as backup while in CA. * Old laptop ready but if it is not inmediately needed I rather have it as backup while in CA.
Topic revision: r14 - 26 May 2020, CarlHolmberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback